OpenStudy (anonymous):

A medical student studying the growth of bacteria in a certain culture has compiled the following data: number of minutes 0 20 number of bacteria 6000 9000 Use these data to find an exponential function of the form Q(t) = Q0e^kt expressing the number of bacteria in the culture as a function of time.How many bacteria are present after 1 hour?

OpenStudy (anonymous):

like finding the equation of a line given two points with a slight difference

OpenStudy (anonymous):

you know it is going to look like \[Q(t)=Q_0\times e^{kt}\] and you have \(Q_0=6000\)

OpenStudy (anonymous):

therefore it is \[Q(t)=6000e^{kt}\] and all you need is \(k\)

OpenStudy (anonymous):

you know if \(t=20\) then \(Q(t)=Q(20)=9000=6000e^{20k}\) set \[9000=6000e^{20k}\] and solve for \(k\)

OpenStudy (anonymous):

first step is to divide by 6000 to get \[1.5=e^{20k}\] then \[\ln(1.5)=20k\] so \[k=\frac{\ln(1.5)}{20}\]

OpenStudy (anonymous):

Q(t)=6000e^(0.203)(t) so this is my equation?